OpenStudy (anonymous):

Solve the equation. Show your work. 16^3 = 4^x This is a fill in the answer

OpenStudy (anonymous):

Well, 16^3 = 16 x 16 x 16 (3 times) and 4^x is 4 times unknown. so, 16 x 16 is, 256,. 256 x 16 = 4096

OpenStudy (anonymous):

then now it's: 4096=4^x now you have to find out how many times 4 times itself into 4096

OpenStudy (anonymous):

4^6

OpenStudy (anonymous):

so x = 6

OpenStudy (anonymous):

16^3 = (4^2)^3 = 4^6 => x = 6